if the point(2,6) is on the graph of y=f(x) what is the value of f(2)? explain
1
Expert's answer
2020-12-01T09:32:08-0500

"f(2)=6"

The graph of a function "f" is the set of points which satisfy the equation"y=f(x)" . That is the point "(x,y)" is on the graph of "f" if and only if "y=f(x)"

Since the graph of "f" is the graph of the equation"y=f(x)" , "f(2)" is the "y" coordinate of the point which corresponds to "(x=2)" . Since the point"(2,6)" is on the graph, then we have "f(2)=6"